TIETOJENKÄSITTELYTIETEEN LAITOS  DEPARTMENT OF COMPUTER SCIENCE
PL 68 (Gustaf Hällströmin katu 2 b)  P.O. Box 68 (Gustaf Hällströmin katu 2 b)
00014 Helsingin yliopisto  FIN-00014 University of Helsinki, Finland
JULKAISUSARJA A  SERIES OF PUBLICATIONS A

\underline{Reports may be ordered from}: Kumpula Science Library, P.O. Box 64, FIN-00014 University of Helsinki, Finland. A-2004-1 M. Koivisto: Sum-product algorithms for the analysis of genetic risks. 155 pp. (Ph.D.

Thesis)

A-2004-2 A. Gurtov: Efficient data transport in wireless overlay networks. 141 pp. (Ph.D.

Thesis)

A-2004-3 K. Vasko: Computational methods and models for paleoecology. 176 pp. (Ph.D.

Thesis)

A-2004-4 P. Sevon: Algorithms for Association-Based Gene Mapping. 101 pp. (Ph.D. Thesis) A-2004-5 J. Viljamaa: Applying Formal Concept Analysis to Extract Framework Reuse Interface

Specifications from Source Code. 206 pp. (Ph.D. Thesis)

A-2004-6 J. Ravantti: Computational Methods for Reconstructing Macromolecular Complexes

from Cryo-Electron Microscopy Images. 100 pp. (Ph.D. Thesis)

A-2004-7 M. Kääriäinen: Learning Small Trees and Graphs that Generalize. 45+49 pp. (Ph.D.

Thesis)

A-2004-8 T. Kivioja: Computational Tools for a Novel Transcriptional Profiling Method. 98 pp.

(Ph.D. Thesis)

A-2004-9 H. Tamm: On Minimality and Size Reduction of One-Tape and Multitape Finite Au-

tomata. 80 pp. (Ph.D. Thesis)

A-2005-1 T. Mielikäinen: Summarization Techniques for Pattern Collections in Data Mining.

201 pp. (Ph.D. Thesis)

A-2005-2 A. Doucet: Advanced Document Description, a Sequential Approach. 161 pp. (Ph.D.

Thesis)

A-2006-1 A. Viljamaa: Specifying Reuse Interfaces for Task-Oriented Framework Specialization.

285 pp. (Ph.D. Thesis)

A-2006-2 S. Tarkoma: Efficient Content-based Routing, Mobility-aware Topologies, and Tempo-

ral Subspace Matching. 198 pp. (Ph.D. Thesis)

A-2006-3 M. Lehtonen: Indexing Heterogeneous XML for Full-Text Search. 185+3 pp. (Ph.D.

Thesis)

A-2006-4 A. Rantanen: Algorithms for ^{13}C Metabolic Flux Analysis. 92+73 pp. (Ph.D. Thesis) A-2006-5 E. Terzi: Problems and Algorithms for Sequence Segmentations. 141 pp. (Ph.D.

Thesis)

A-2007-1 P. Sarolahti: TCP Performance in Heterogeneous Wireless Networks. (Ph.D. Thesis) A-2007-2 M. Raento: Exploring privacy for ubiquitous computing: Tools, methods and experi-

ments. (Ph.D. Thesis)

A-2007-3 L. Aunimo: Methods for Answer Extraction in Textual Question Answering. 127+18

pp. (Ph.D. Thesis)

A-2007-4 T. Roos: Statistical and Information-Theoretic Methods for Data Analysis. 82+75 pp. (Ph.D. Thesis)

A-2007-5 S. Leggio: A Decentralized Session Management Framework for Heterogeneous Ad-Hoc

and Fixed Networks. 230 pp. (Ph.D. Thesis)

A-2007-6 O. Riva: Middleware for Mobile Sensing Applications in Urban Environments. 195 pp.

(Ph.D. Thesis)

A-2007-7 K. Palin: Computational Methods for Locating and Analyzing Conserved Gene Regu-

latory DNA Elements. 130 pp. (Ph.D. Thesis)

A-2008-1 I. Autio: Modeling Efficient Classification as a Process of Confidence Assessment and

Delegation. 212 pp. (Ph.D. Thesis)

A-2008-2 J. Kangasharju: XML Messaging for Mobile Devices. 24+255 pp. (Ph.D. Thesis). A-2008-3 N. Haiminen: Mining Sequential Data – in Search of Segmental Structures. 60+78 pp.

(Ph.D. Thesis)

A-2008-4 J. Korhonen: IP Mobility in Wireless Operator Networks. (Ph.D. Thesis) A-2008-5 J.T. Lindgren: Learning nonlinear visual processing from natural images. 100+64 pp.

(Ph.D. Thesis)

A-2009-1 K. Hätönen: Data mining for telecommunications network log analysis. 153 pp. (Ph.D.

Thesis)

A-2009-2 T. Silander: The Most Probable Bayesian Network and Beyond. (Ph.D. Thesis) A-2009-3 K. Laasonen: Mining Cell Transition Data. 148 pp. (Ph.D. Thesis) A-2009-4 P. Miettinen: Matrix Decomposition Methods for Data Mining: Computational Com-

plexity and Algorithms. 164+6 pp. (Ph.D. Thesis)

A-2009-5 J. Suomela: Optimisation Problems in Wireless Sensor Networks: Local Algorithms

and Local Graphs. 106+96 pp. (Ph.D. Thesis)

A-2009-6 U. Köster: A Probabilistic Approach to the Primary Visual Cortex. 168 pp. (Ph.D.

Thesis)

A-2009-7 P. Nurmi: Identifying Meaningful Places. 83 pp. (Ph.D. Thesis) A-2009-8 J. Makkonen: Semantic Classes in Topic Detection and Tracking. 155 pp. (Ph.D.

Thesis)

A-2009-9 P. Rastas: Computational Techniques for Haplotype Inference and for Local Alignment

Significance. 64+50 pp. (Ph.D. Thesis)

A-2009-10 T. Mononen: Computing the Stochastic Complexity of Simple Probabilistic Graphical

Models. 60+46 pp. (Ph.D. Thesis)

A-2009-11 P. Kontkanen: Computationally Effcient Methods for MDL-Optimal Density Estima-

tion and Data Clustering. 75+64 pp. (Ph.D. Thesis)

A-2010-1 M. Lukk: Construction of a global map of human gene expression - the process, tools

and analysis. 120 pp. (Ph.D. Thesis)

A-2010-2 W. Hämäläinen: Efficient search for statistically significant dependency rules in binary data. (Ph.D. Thesis)